# Sample-efficient Imitative Multi-token Decision Transformer for Generalizable Real World Driving

### v1

@@ 4.1 Experimental Setup
Agent Action Space Sim Agent Off-Road Rate (%) Collision Rate (%) Kinematic Infeasibility (%) ADE (m) Route Progress Ratio (%)
@@
Wayformer[37] Delta - 7.89 10.68 5.40 2.38 123.58
BC[38] Delta - 4.14$\pm$2.04 5.83$\pm$1.09 0.18$\pm$0.16 6.28$\pm$1.93 79.58$\pm$24.98
@@
SimDT(ours) Bicycle Playback 3.36$\pm$0.04 2.65$\pm$0.06 0.00$\pm$0.00 6.73$\pm$0.41 -
@@
Table 1: Performance evaluation are done against IDM simulation agents. Agents run without any termination conditions in WOD1.2 evaluation dataset. Action space is continuous unless denoted with D (discrete). Waymax benchmark table is used as we use exact same experiment settings.
Dataset, simulator and metrics. Training and Experiments are done based on Waymo Open Dataset and Waymax simulator. Waymax provides embedded support for reinforcement learning and diverse scenarios drawn from real driving data. Waymax incorporate with Waymo Open Motion Dataset (WOMD) which provides $531,101$ real-world driving scenarios for training and $44,096$ scenarios for validation, each scenario contains 90 frames of data. Specifically, WOMD v1.2 and exact same metrics (off-road rate, collision rate, kinematic infeasibility, average displacement error (ADE)) from Waymax are used to benchmark with the paper.
Implementation Detail. Models of various sizes are developed to quickly conduct ablation studies and assess final performance effectively. Raw observation takes nearest ego vehicle, 15 nearest dynamic obstacles, 250 of closest roadgraph elements, traffic signals and position goal as input. The total size for each step observation is $7050$ and feature extraction is applied to reduce the total size. SimDT(tiny) has 256 tokens for each element of $(s,a,g)$ pair, 6 blocks, 16 attention head and in total 7.7 million parameters. SimDT(small) has 384 tokens for each element of $(s,a,g)$ pair, 10 blocks, 16 attention head and in total 22.2 million parameters. Both models use context length with value 10, which means causal transformer has access to past 10 $(s,a,g)$ pairs.
4.2 Benchmark Comparison
SimDT is evaluated using Intelligent Driving Model (IDM)[40] as the simulated agent. SimDT achieves Off-Road Rate of 3.36%, Collision Rate of 2.65%, Kinematic Infeasibility of 0.00%, and ADE of 6.73m. SimDT significantly outperforms them in collision rate and being second in off-road rate against other learning-based approaches. Compared same reinforcement learning category method, SimDT demonstrates a substantial reduction in Off-Road Rate and Collision Rate than DQN. Suggesting that our method is more effective at keeping the vehicle on the road and avoiding accidents. The Off-Road Rate of SimDT is higher than the best performing BC ’Bicycle (D)’ model by 2%. Similarly, the Collision Rate of SimDT shows a 1.94 percentage point improvement over the same BC model. This improvement in safety-critical metrics highlights the robustness of SimDT in real-world driving scenarios.
When compared to expert demonstrations, SimDT achieves competitive results in terms of safety metrics Collision Rate are within the same magnitude as those reported by the experts. However, the ADE of SimDT is notably higher at 6.73m, which is approximately 6 meters away from the expert models. This suggests that SimDT learns a safe and feasible policy but different from the expert recording. While the ADE for SimDT is higher than that of other imitation learning models, it is important to note that ADE alone may not capture the complete picture of driving performance. The emphasis on safety and kinematic feasibility by SimDT may contribute to a cautious driving style, which can result in a slightly higher ADE but with significantly safer outcomes.
4.3 Ablation Study
Agent Off-Road Rate (%) Collision Rate (%) Kinematic Infeasibility (%) ADE (m)
DT(tiny) 5.86 3.43 0.00$\pm$0.00 7.64
DT(tiny) + PER 4.28 3.27 0.00$\pm$0.00 7.26
DT(tiny) + PER + OPA 3.73 2.79 0.00$\pm$0.00 7.05
DT(tiny) + PER + OPA + 2 token prediction 3.74 2.72 0.00$\pm$0.00 7.09
DT(tiny) + PER + OPA + 3 token prediction 3.57 2.54 0.00$\pm$0.00 6.97
DT(small) + PER + OPA + 3 token prediction 3.36 2.65 0.00$\pm$0.00 6.73
@@ 4.3 Ablation Study
Table 2: Ablation Study. PER is prioritized experience replay, OPA is online-policy adaption.
@@
Figure 3: Illustration of Data Selection for Prioritized Experience Replay: 3(a) is chosen because its uncommon expert behavior that need to slow down while steer to the right to keep lane. 3(b) illustrates a rare parking situation and highlights a case that was picked because it had the most mistakes when looking at the whole series of actions. 3(c) is kept as the suboptimal action taken in that situation was not reproduced given the corresponding low return-to-go.
@@
Prioritized Experience Replay for Decision Transformer. Since our proposed imitative reinforcement learning can obtain almost infinite amount of dataset through online interaction, the ability of prioritized experience replay becomes critical for sample efficiency. Compare to pure Decision Transformer, the model which adapts PER has 1.58% and 0.16% reduction in off-road and collision rate. Decision Transformer with PER is able to reach same performance with 80% of data. There are three types of the data that is preferentially stored for PER (Fig. 3). The initial category encompasses instances wherein a discernible discrepancy arises between the predicted actions of the learning model and those executed by an expert. The second category pertains to scenarios wherein the cumulative action loss associated with a particular trajectory is substantially elevated, a phenomenon that predominantly transpires within the confines of rare encountered environmental conditions. The third category is representative of situations where trajectories indicative of suboptimal online adaptation are documented, highlighting the model’s challenges in identifying and rectifying suboptimal behaviors. The sample-efficient leanrnig curve can be found in Appendix. 6
Multi-token Decision Transformer. Due to physics limitation of real world vehicles such as inertia and momentum, actions taken at current time-step can significantly affect the following time-step actions. Current state has effect to near future actions steps. eg. reckless pedestrian crossing can cause emergency breaking for ego vehicle and it takes at least few steps to finish. It is important for Neural Network to understand return-to-go and sequence action consequence during the training stage. On the other hand, Focusing only on predicting one token at a time makes a model too sensitive to the immediate context, overlooking the need to consider longer sequences of token for better understanding. Multi-token prediction allows for a more nuanced grasp of world interaction with less data. Compared with single-token prediction, 3-token SimDT has 0.16% and 0.25% improvement in off-road rate and collision rate.
@@
Figure 4: Attention map comparison for single-token and multi-token prediction. Multi-token prediction network has more diverse attention field.
@@
5 Conclusion and Discussion
We introduces SmiDT, an innovative approach to sequence modeling based reinforcement learning, particularly targeted for the complexities of real-world driving scenarios. Our fully online imitative Decision Transformer pipeline is adept at handling diverse data distributions found within extensive driving datasets, ensuring wide applicability and robustness. By implementing a multi-token Decision Transformer that integrates receding horizon control, we improve the model’s ability to predict over longer horizons and extend its attention span across broader contexts. Furthermore, the incorporation of prioritized experience replay within our framework enhances the sample efficiency of training, allowing for more effective learning from large-scale datasets. Our work can also benefit other real-world robotics tasks that demand sample-efficient imitative reinforcement learning.
Limitation. Due to computational constraints, we couldn’t train a larger network with increased embedding sizes, more transformer blocks, additional attention heads, and extended context length. A longer context would enhance the model’s grasp of its environment, potentially improving its capability for both high-level task planning and low-level action planning.
Acknowledgments
If a paper is accepted, the final camera-ready version will (and probably should) include acknowledgments. All acknowledgments go at the end of the paper, including thanks to reviewers who gave useful comments, to colleagues who contributed to the ideas, and to funding agencies and corporate sponsors that provided financial support.
Appendix
A. Learning cure
where model 1 is Pure Decision transformer, and model 2 has Decision Transformer with Prioritized Experience Replay. Figure shows our Sample-efficient Imitative Pipeline converges faster and has better performance.
Figure 6: Learning Curve. where Model 1 is Pure Decision transformer, and Model 2 has Decision Transformer with Prioritized Experience Replay. Figure shows our Sample-efficient Imitative Pipeline converges faster and has better performance.
B. Metrics Definition
Collision rate This metric checks for overlap between bounding boxes of objects in a 2D top-down view at the same time step to determine if a collision has occurred.
Off-Road rate indicates the percentage whether the vehicle is driving within the road boundaries, with any deviation to the right of the road’s edge considered off-road.
Kinematic Infeasibility Metric is binary metric assesses whether a vehicle’s transition between two consecutive states is within predefined acceleration and steering curvature limits, based on inverse kinematics.
Average Displacement Error (ADE) calculates the mean L2 distance between the vehicle’s simulated position and its logged position at corresponding time steps across the entire trajectory.
Route Progress Ratio calculates the proportion of the planned route completed by the vehicle, based on the closest point along the path at a given time step. Route Progress Ratio feature is not released yet and benchmark in this paper will skip this metric.

### v2

@@ 4.1 Experimental Setup
Dataset, simulator and metrics. Training and experiments are conducted using the Waymo Open Dataset and the Waymax simulator. Waymax provides embedded support for reinforcement learning and diverse scenarios drawn from real driving data. It integrates with the Waymo Open Motion Dataset (WOMD), which offers 531,101 real-world driving scenarios for training and 44,096 scenarios for validation. Each scenario contains 90 frames of data. Specifically, WOMD v1.2 and the exact same metrics (off-road rate, collision rate, kinematic infeasibility, average displacement error (ADE)) from Waymax are used for benchmarking with the paper. Evaluations are conducted in both open-loop and closed-loop settings for a wider range of performance comparisons.
Implementation Detail. Models of various sizes are developed to quickly conduct ablation studies and assess final performance effectively. Raw observation takes ego vehicle, 15 nearest dynamic obstacles, 300 of closest roadgraph elements, traffic signals and position goal as input. The total size for each step observation is $8892$ and feature extraction is applied to reduce the total size. SimDT(small) has 384 tokens for each element of $(s,a,g)$ pair, 10 blocks, 16 attention heads and in total 22.6 million parameters. SimDT(small) has 512 tokens for each element of $(s,a,g)$ pair, 15 blocks, 16 attention heads and in total 53.2 million parameters. Both models use context length with value 10, which means the causal transformer has access to past 10 $(s,a,g)$ pairs.
4.2 Benchmark Comparison
SimDT is evaluated both in open-loop and closed-loop settings for comprehensive comparison with various algorithms. During closed-loop evaluation, Intelligent Driving Model (IDM)(Treiber et al., 2000) is deployed as the simulated agent. SimDT achieves an Off-Road Rate of 3.36%, Collision Rate of 2.65%, Kinematic Infeasibility of 0.00%, and ADE of 6.73m. SimDT significantly outperforms them in collision rate and is second in off-road rate against other learning-based approaches. Compared with the same RL category method, SimDT demonstrates a substantial reduction in Off-Road Rate and Collision Rate than DQN by 45.2%. Similarly, the Collision Rate of SimDT shows 41% improvement over the Behavior Cloning (BC) model. Suggesting that our method is more effective at keeping the vehicle on the road and avoiding accidents. This improvement in safety-critical metrics highlights the robustness of SimDT in real-world driving scenarios.
The open-loop evaluation (Figure 2) shows SimDT has comparable results with the state-of-the-art algorithm BC-SAC. BC again shows a solid performance with a low failure rate of 4.35% and a high route progress ratio of 99.00%. BC-SAC improves upon SAC, reduces the failure rate to 3.35% and increases route progress to 95.26%. Remarkably, our proposed SimDT method achieves a moderate failure rate of 3.87% while surpassing others with a route progress ratio of 105.63%, indicating not only effective navigation but also the potential for discovering more efficient routes. These results underline the importance of selecting appropriate training simulations and methods for optimizing autonomous navigation systems.
When compared to expert demonstrations, SimDT achieves competitive results in terms of safety metrics. Collision Rates are within the same magnitude as those reported by the experts. However, the ADE of SimDT is notably higher at 6.73m, which is approximately 6 meters away from the expert models. This suggests that SimDT learns a safe and feasible policy but is different from the expert recording. While the ADE for SimDT is higher than that of other imitation learning models, it is important to note that ADE alone may not capture the complete picture of driving performance. The emphasis on safety and kinematic feasibility by SimDT may contribute to a cautious driving style, which can result in a slightly higher ADE but with significantly safer outcomes.
Agent Action Space Train Sim Agent Off-Road Rate (%) Collision Rate (%) Kinematic Infeasibility (%) ADE (m) Route Progress Ratio (%)
@@ 4.2 Benchmark Comparison
Wayformer (Nayakanti et al., 2022) Delta - 7.89 10.68 5.40 2.38 123.58
BC(Argall et al., 2009) Delta - 4.14$\pm$2.04 5.83$\pm$1.09 0.18$\pm$0.16 6.28$\pm$1.93 79.58$\pm$24.98
@@
SimDT(ours) Bicycle Playback 3.52$\pm$0.26 2.69$\pm$0.10 0.00$\pm$0.00 7.14$\pm$0.63 106.47$\pm$2.8
@@
Table 1: Closed-loop Benchmark. Performance evaluations are done against IDM simulation agents. Agents run without any termination conditions. Models report mean and standard deviation over 3 seeds. Action space is continuous unless denoted with D (discrete).
Method Train Sim Agent Failure Rate (%) Route Progress Ratio (%)
BC(Argall et al., 2009) - 4.35$\pm$0.27 99.00$\pm$0.39
SAC(Haarnoja et al., 2018) Playback 6.66$\pm$0.44 77.82$\pm$8.21
BC-SAC(Lu et al., 2023) Playback 3.35$\pm$0.31 95.26$\pm$8.64
SimDT (ours) Playback 3.87$\pm$0.26 105.63$\pm$2.31
Table 2: Open-loop Benchmark. Failure rates (lower is better) of BC-SAC and baselines on different training/evaluation subsets. The failure rate is the percentage of the run segments that have at least one collision or off-road event. Models report mean and standard deviation over 3 seeds.
4.3 Ablation Study and Analysis
Agent Off-Road Rate (%) Collision Rate (%) Kinematic Infeasibility ADE (m)
DT(small) + 1 token prediction 6.21$\pm$0.35 3.62$\pm$0.18 0.00$\pm$0.00 8.32$\pm$0.68
DT(small) + PER + 1 token prediction 4.59$\pm$0.31 3.42$\pm$0.15 0.00$\pm$0.00 7.95$\pm$0.69
DT(small) + PER + OPA + 1 token prediction 3.97$\pm$0.24 2.92$\pm$0.13 0.00$\pm$0.00 7.64$\pm$0.59
DT(small) + PER + OPA + 3 token prediction 3.82$\pm$0.25 2.65$\pm$0.11 0.00$\pm$0.00 7.52$\pm$0.55
DT(small) + PER + OPA + 5 token prediction 3.75$\pm$0.25 2.59$\pm$0.10 0.00$\pm$0.00 7.45$\pm$0.57
DT(small) + PER + OPA + 7 token prediction 4.14$\pm$0.30 3.33$\pm$0.11 0.00$\pm$0.00 8.19$\pm$0.59
DT(median) + PER + OPA + 3 token prediction 3.52$\pm$0.26 2.69$\pm$0.10 0.00$\pm$0.00 7.14$\pm$0.63
@@ 4.3 Ablation Study and Analysis
Table 3: Ablation Study. PER is prioritized experience replay, OPA is online policy adaptation.
Figure 3: Comparison of learning curves for DT(small) settings with and without Prioritized Experience Replay. The metric for reward indicates that higher values are better, whereas for off-road incidents, overlap, and log divergence, lower values are better. Models report mean and standard deviation over 3 seeds.
@@
Figure 4: Illustration of Data Selection for Prioritized Experience Replay: 4(a) is chosen because of its uncommon expert behavior that needs to slow down while steering to the right to keep lane. 4(b) illustrates a rare parking situation and was picked because it had the most mistakes when looking at the whole series of actions. 4(c) is kept as the suboptimal (collision) action taken in that situation was not reproduced given the corresponding low return-to-go.
@@
Prioritized Experience Replay for Decision Transformer. Compared to the baseline Decision Transformer model, integrating Prioritized Experience Replay results in a significant reduction of off-road incidents and collision rates by 26.1% and 5.5%, respectively (Table 3). Moreover, the enhanced Decision Transformer model achieves comparable performance using only 60% of the scenarios. The comparison of learning curves with and without Prioritized Experience Replay is presented in Figure 3. Multiple experiments are conducted with different initial seeds to ensure reproducibility. Three types of data are preferentially stored for PER (Figure 4). The initial category encompasses instances wherein a discernible discrepancy arises between the predicted actions of the learning model and those executed by an expert. The second category pertains to scenarios wherein the cumulative action loss associated with a particular trajectory is substantially elevated, a phenomenon that predominantly transpires within the confines of rare encountered environmental conditions. The third category is representative of situations where trajectories indicative of suboptimal online adaptation are documented, highlighting the model’s challenges in identifying and rectifying suboptimal behaviors.
A key aspect of the Decision Transformer is its ability to model the relationship between state, action, and return-to-go ($g_{t}$). Given the same state observation, expert actions are associated with a higher $g_{t}$, and sub-optimal actions with a lower $g_{t}$. Ideally, DT should learn to act differently (e.g. choose an expert action, avoid driving out of the lane, or prevent collisions) based on the $g_{t}$ value. When a predicted action significantly diverges from the ground truth action given state observation and specific $g_{t}$, it indicates a misinterpretation of the environment. Intuitively, such misbehaved experiences need further reinforcement through the replay buffer.
From a high-level perspective, training on large-scale real-world data presents uneven scenario distributions. When the network begins to generalize common scenarios, rare scenarios like Figures 4(a) and 4(b) are undertrained. At this point, the network performs relatively well on common scenarios (low action loss) but not on rare scenarios (high action loss). Consequently, the replay buffer starts storing more of these rare scenarios. Criterion 2 is applied with accumulated action loss to prioritize such scenarios. Figure 4(c) illustrates a policy rollout during online adaptation where suboptimal actions cause changes in $g_{t}$. What is important is the record of the corresponding $g_{t}$ with respect to action, where expert actions and sub-optimal actions are respectively given positive and negative rewards. eg. Figure 4(c) shows the ego vehicle driving out of the lane with a recorded decrease in $g_{t}$ value -2. It is crucial for the neural network to understand which actions cause a -2 value drop in $g_{t}$. These suboptimal state-action-return sequences are also important as expert driving data doesn’t contain such behavior. The next time the agent encounters a similar scenario, the correct $g_{t}$ will help prevent out-of-lane actions.
Multi-token Decision Transformer. In ablation study detailed in Table 3, we compare the performance of single-token prediction against multi-token predictions. The 3-token and 5-token SimDT variants demonstrate improvements of 3.78% and 6.80% in off-road, 8.56% and 11.3% in collision rates respectively. This improvement underscores the importance of considering future actions, influenced by real-world vehicle dynamics like inertia and momentum, which single-token models may overlook due to their focus on immediate contexts. The attention map in Figure 5 shows that multi-token prediction has wider attention fields towards long-term context. However, extending predictions to 7 tokens resulted in slight performance degradation, suggesting existing difficulties in trajectory stitching for larger sequences modeling by Decision Transformer. These results indicate that while multi-token prediction models offer a more nuanced understanding of environmental interactions, the optimization of larger token prediction poses a challenge that needs further study.
@@
Figure 5: Attention map comparison for single-token and multi-token prediction. Multi-token prediction network has a more diverse attention field. It indicates that multi-token prediction is less susceptible to immediate contextual patterns.
5 Conclusion
We present SmiDT, a novel sequence modeling-based reinforcement learning approach for closed-loop driving and complex real-world driving scenarios. Our fully online imitative Decision Transformer pipeline is adept at handling diverse data distributions found within extensive driving datasets, ensuring wide applicability and robustness. By implementing a multi-token Decision Transformer that integrates receding horizon control, we improve the model’s ability to predict over longer horizons and extend its attention span across broader contexts. Furthermore, the incorporation of prioritized experience replay within our framework enhances the sample efficiency of training, allowing for more effective learning from large-scale datasets. Our work can also benefit other real-world robotics tasks that demand sample-efficient learning from expert demonstration and online adaptation.
Appendix
Figure 6: The general training procedure of SimDT. Real-world driving scenarios are reproduced in the simulator and the produced transitions $(s,a,r)$ are recorded to replay buffer. HindsightReturnRelabeling is performed to transform transitions to state-action-return pairs $(s,a,g)$. During training, prioritized experience replay is performed for sample efficiency. Finally, online adaptation is used for better performance.
A. Metrics
The Off-road Rate indicates the percentage of whether the vehicle is driving within the road boundaries, with any deviation to the right of the road’s edge considered off-road.
The Kinematic Infeasibility Metric is a binary metric that assesses whether a vehicle’s transition between two consecutive states is within predefined acceleration and steering curvature limits, based on inverse kinematics.
Average Displacement Error (ADE) calculates the mean euclidean distance between the vehicle’s simulated position and its logged position at corresponding time steps across the entire trajectory.
Route Progress Ratio calculates the proportion of the planned route completed by the vehicle, based on the closest point along the path at a given time step. The Route Progress Ratio feature is not released yet and the benchmark in this paper will skip this metric.
B. Choice for imitation reward log_divergence threshold
The log_divergence is the distance between the logged position and the actual agent position. The key here is to reward the system when the agent is close to expert human actions. When performing real-world human demonstration through inverse kinematics, the log_divergence is close to zero. The reason to set it to 0.2 is to reward policy agents. However, we consider 0.2 to be the maximum that we can set as values bigger than 0.2 can lead to a collision and out-of-lane actions which should not be rewarded. In general, we do not see a significant difference when varying values from 0.05 to 0.2. The reason can be 1. Change of threshold values has no impact on reward functions during the offline reinforcement learning phase. 2. Values under 0.2 are still relatively small for agents to follow during the online exploration phase.
C. Comparison with BC-SAC
Despite SimDT and BC-SAC both trying to combine imitation learning and reinforcement learning, they use different approaches to achieve this goal. The key for BC-SAC is to combine reinforcement learning cost with imitation learning cost during training. Our proposed imitative reinforcement learning pipeline is a reinforcement learning method that combines rewards for imitating expert demonstration.
D. Coefficient for Multi-token Training
In the main body of the text, the modified loss function $L_{ma}$ is designed to enhance the network’s learning capability regarding current step action predictions through the strategic use of coefficients $\alpha$, $\beta$, …, and $\gamma$. The values that are adapted during training follow a decay setting as we think the near future token prediction is more important than further future tokens.
Method $\alpha$ $\beta$ $\gamma$ $\delta$ $\sigma$ $\phi$ $\omega$
SimDT (1-token) 1.00 - - - - - -
SimDT (3-token) 0.80 0.40 0.20 - - - -
SimDT (5-token) 0.80 0.30 0.15 0.10 0.08 - -
SimDT (7-token) 0.80 0.25 0.12 0.10 0.08 0.06 0.05
Figure 7: The general training procedure of SimDT. We minimize the action loss over a horizon to thereby minimize the deviation of the resultant predicted trajectory from the reference trajectory.